# Bayesian Inference General Procedures for A Single-subject Test Study

### v2

@@ 3.4 Comparison Study with Other Frameworks
This section is dedicated to assessing the performance of the proposed BIGPAST methodology and existing approaches when a control group is present. The first approach is a variant of BIGPAST based on the maximum a posteriori (MAP) estimation. Specifically, procedures 3-5 of the BIGPAST are simplified into the following two steps:
  1. 3
  2. 4
    Given the significance level $\alpha^{*}$, the credible interval of $x^{*}$ given $\mathbf{x}$ is $[x^{\mathrm{MAP}}_{\alpha^{*}/2},x^{\mathrm{MAP}}_{1-\alpha^{*}/2}]$, where $x^{\mathrm{MAP}}_{\alpha^{*}/2}$ and $x^{\mathrm{MAP}}_{1-\alpha^{*}/2}$ are the $\alpha^{*}/2$ and $21-\alpha^{*}/$ 2 quantiles of $\boldsymbol{\pi}(\hat{\alpha}_{\mathrm{MAP}},\hat{\nu}_{\mathrm{MAP}},\hat{\xi}_{\mathrm{MAP}},\hat{\omega}_{\mathrm{MAP}}|\mathbf{x})$, respectively. If $x^{*}$ does not fall within this credible interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\alpha^{*}$.
The second approach is a non-Bayesian approach that relies on maximum likelihood estimation (MLE). This approach operates under the sole assumption that the control group follows a skewed Student’s $t$ distribution. The MLE procedures can be outlined as follows:
@@
    Given the significance level $\alpha^{*}$, the confidence interval of $x^{*}$ given $\mathbf{x}$ is $[x^{\mathrm{MLE}}_{\alpha^{*}/2},x^{\mathrm{MLE}}_{1-\alpha^{*}/2}]$, where $x^{\mathrm{MLE}}_{\alpha^{*}/2}$ and $x^{\mathrm{MLE}}_{1-\alpha^{*}/2}$ are the $\alpha^{*}/2$ and $21-\alpha^{*}/$ 2 quantiles of $f_{c}(\hat{\alpha}_{\mathrm{MLE}},\hat{\nu}_{\mathrm{MLE}},\hat{\xi}_{\mathrm{MLE}},\hat{\omega}_{\mathrm{MLE}})$, respectively. If $x^{*}$ does not fall within this confidence interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\alpha^{*}$.
The third approach is a non-parametric (NP) approach. Given group data $\mathbf{x}$ and a significance level $\alpha^{*}$, we can derive the empirical confidence interval $[x^{\mathrm{NP}}_{(\lfloor n*\alpha^{*}/2\rfloor)},x^{\mathrm{NP}}_{(\lceil n*(1-\alpha^{*}/2)\rceil)}]$ where $x^{\mathrm{NP}}_{(\lfloor n*\alpha^{*}/2\rfloor)}$ and $x^{\mathrm{NP}}_{(\lceil n*(1-\alpha^{*}/2)\rceil)}$ are the empirical quantiles of $\mathbf{x}$ at $\alpha^{*}/2$ and $21-\alpha^{*}/$ 2, respectively. If $x^{*}$ does not fall within this empirical confidence interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\alpha^{*}$. The NP approach does not rely on any distributional assumptions. Given that our BIGPAST methodology employs the Metropolis-Hastings sampler in Step 3, we will refer to the aforementioned four approaches: MH, MAP, MLE, and NP, respectively.
@@
$\mathrm{FDR}_{i}^{\mathcal{M}}\coloneqq\frac{\mathrm{FP}_{i}^{\mathcal{M}}}{\mathrm{TP}_{i}^{\mathcal{M}}+\mathrm{FP}_{i}^{\mathcal{M}}},\quad\mathrm{ACC}_{i}^{\mathcal{M}}\coloneqq\frac{\mathrm{TP}_{i}^{\mathcal{M}}+\mathrm{TN}_{i}^{\mathcal{M}}}{K},\quad i=1,2,\ldots,N,$
where $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{TN}$ represent the number of true positive, false positive and true negative, respectively. $\mathcal{M}$ could be one of MH, MAP, MLE, NP.
The parameters settings are as follows: we consider ten combinations of $(\alpha,\nu)$: (0,3), (1,3), (3,3), (5,5), (10,10), (20,20), (30,30), (50,50), (5,50) and (50,5) with fixed $\xi=-2$ and $\omega=2$. In MH approach, we let $m=5000$, $s=100$ and $c_{\nu}=1$. The other parameter settings are $N=400$, $n=100$, $k_{1}=k_{2}=1000$, $\alpha^{*}=0.05$. The results are presented in Table 5 and Figure 4.
Table 5 displays the FDR and ACC of the four approaches. The MH and MAP approaches consistently outperform the NP and MLE approaches in terms of FDR and ACC. When ${\left\lvert\alpha\right\rvert}\leq 3$ and $\nu\leq 3$, both the FDR and ACC of the MAP approach surpass those of the MH approach. This is expected, as the MAP approach provides more accurate estimation of the underlying parameters of the skewed Student’s $t$ distribution when ${\left\lvert\alpha\right\rvert}$ and $\nu$ are small, as is shown in column 3 of Table 2. When ${\left\lvert\alpha\right\rvert}$ or $\nu$ is large, the FDR and ACC of the MH approach perform better than those of the MAP approach. This can be attributed to the randomness of parameters in the Metropolis-Hastings sampler. However, the MAP approach fails to estimate the underlying parameters when $\alpha$ or $\nu$ is large, as is shown in column 4 of Table 2, and lacks randomness of parameters. The boxplots in Figure 4 clearly illustrate the comparison between the MH and MAP approaches. The results illustrate that the FDR of MH (i.e. BIGPAST) is generally lower than others, and the ACC of MH (i.e. BIGPAST) is generally greater than other methods in conducting hypothesis tests for single-subject data.
@@
Table 5: The average false discovery rate (FDR) and accuracy (ACC) for the MH, MAP, MLE, and NP approaches, calculated over 400 repetitions. In this context, MH refers to the original BIGPAST method, MAP denotes the Bayesian inference framework based on maximum a posteriori estimation, MLE represents the method based on maximum likelihood estimation, and NP signifies the non-parametric method.
@@ 4 Real Data Analysis
This dataset is derived from a mild traumatic brain injury (mTBI) study conducted by Innovision IP Ltd. It comprises functional time series of Magnetoencephalography (MEG) data and an anatomical T1-weighted Magnetic Resonance Imaging (MRI) scan for 103 healthy subjects and one patient with mild traumatic brain injury. The MEG data, recorded from 306 sensors, was pre-processed using the ‘MNE-Python’ package (Gramfort et al., 2013). The MEG data was analysed to reconstruct the source generators using a minimum norm inverse method. The MRI data was pre-processed with the ‘FreeSurfer’ package (Han et al., 2006). The standard FreeSurfer pipeline (detailed on their website) was followed and includes motion correction, intensity normalisation, skull-stripping, normalisation to template space and segmentation of the cortex and subcortical structures using the Desikan-Killiany atlas. This results in 68 regions covering the entire cortex. For each cortical region, we calculate the average of source powers within this region. This process yields a control group dataset with dimensions $103\times 68\times 100$ and a patient dataset with dimensions $68\times 100$. To simplify, we treat the epochs as repeated observations and take average over them to reduce the data’s complexity. Consequently, the control group dataset has dimensions $103\times 68$, and the patient dataset is reduced to a vector of length $68$.
@@
For each brain region of the Desikan-Killiany cortical atlas, we apply the BIGPAST method to the control group and patient dataset to test whether the patient significantly differs from the control group. In this real data analysis, we know that the single-subject has mTBI. However, in practice, we can calculate the observation of a new client and conduct the hypothesis test using BIGPAST to determine whether the client significantly differs from the control group.
We further explore the impact of waveband frequency on the comparison between a control group and a single-subject. The wavebands are arranged in ascending order of frequency: delta (1-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (12-30 Hz), and gamma (30-45 Hz). The subsequent results also include their findings, denoted as CG for brevity. The prespecified significance level is 0.05.
(CG, BIGPAST) delta theta alpha beta gamma
@@
Table 6: The test results of the Crawford-Garthwaite (CG) and BIGPAST approaches, both based on the normal and skewed Student’s $t$ assumptions. The indices corresponding to specific brain cortical regions are 1: lh-caudal anterior cingulate, 2: lh-caudal middle frontal, 4: lh-entorhinal, 6: lh-fusiform, 10: lh-isthums cingulate, 12: lh-lateral orbitofrontal, 14: lh-medial orbitofrontal, 17: lh-parahippocampal, 22: lh-postcentral, 24: lh-precentral, 26: lh-rostral anterior cingulate, 35: rh-caudal anterior cingulate, 38: rh-entorhinal, 40: rh-fusiform, 47: rh-lingual, 50: rh-paracentral, 51: rh-parahippocampal, 58: rh-precentral, 59: rh-precuneus, 62: rh-superior frontal. The prefixes ‘lh-’ and ‘rh-’ indicate the left and right hemispheres of the brain, respectively. ‘NA’ stands for ‘not available’, and ‘Others’ represents the remaining brain cortical indices not listed in the column, but can be found the indices in Table 10.
@@
Figure 5: The comparison results between the Crawford-Garthwaite (CG) and BIGPAST approaches in terms of density estimations. The light-red lines represent the skewed Student’s $t$ densities, with parameters drawn from the BIGPAST approach’s posterior distribution. The light green lines depict the normal densities, with parameters drawn from the CG approach’s posterior distribution. The CG test is based on 10,000 densities, while the BIGPAST test is based on 5,000 densities (the first 5,000 densities are discarded for burn-in). For clarity, each panel only displays 200 densities drawn from the CG and BIGPAST posterior distribution. The red vertical line on the $x$ axis represents the observation of a single-subject (i.e., a patient with mTBI). A histogram is used to demonstrate the goodness of fit for both the CG and BIGPAST approaches.
In this real data analysis, the null hypothesis is two-sided, with one heavy tail and one light tail in the skewed Student’s $t$ distribution. The cortical region indices in row 3 of Table 6 exhibit a positive skewness parameter, i.e., $\alpha>0$. Consequently, the light tail is on the left, and the heavy tail is on the right. For all cortical region indices in row 3 of Table 6, the single-subject observation is located at the light tail of the skewed Student’s $t$ distribution. This positioning suggests a high probability of rejecting the null hypothesis based on the skewed Student’s $t$ assumption, as illustrated in Figure 5(a). Similarly, for cortical regions 6 (lh-fusion, gamma band), 14 (lh-medial orbitofrontal,theta and beta bands), 40 (rh-fusiform, theta and gamma bands), and 47 (rh-lingual, beta band), the single-subject observation is located at the heavy tail of the skewed Student’s $t$ distribution, as shown in Figure 5(b). The skewed Student’s $t$ assumption is more appropriate for these cortical regions than the assumption of normal distribution. The BIGPAST approach outperforms the CG approach in detecting the single-subject observation at the heavy tail of the skewed Student’s $t$ distribution, as demonstrated in Figure 5. Because the BIGPAST approach corrects the model misspecification error at both tails compared to the CG approach. By combining rows 2 and 3 of Table 6 and Figure 5, we conclude that the BIGPAST approach is more reliable than the CG approach in hypothesis tests for single-subject data when the control group data exhibits skewness.
@@ 5 Conclusion
In this paper, we proposed a novel Bayesian inference framework for the abnormality detection on a single-subject versus a control group. Under the assumption of a normal distribution, the $t$-score (Crawford and Howell, 1998) is capable of managing scenarios where the sample size is fewer than 30 (Machin et al., 2018). However, in neuroscience, the number of epochs often exceeds 100, sometimes even reaching 1000, and the distributions—whether over epochs or subjects—exhibit skewness. Under these conditions, methods predicated on the assumption of normality cease to be effective. Our proposed methodology, BIGPAST, leverages the skewed Student’s $t$ distribution, effectively capturing the inherent asymmetry of the data.
In terms of accuracy, power, and Type I error, our proposed BIGPAST methodology outperforms the existing Crawford-Garthwaite (CG) approach. Particularly when the control group data exhibits skewness, BIGPAST provides a more reliable solution for hypothesis testing on single-subject data. Furthermore, BIGPAST demonstrates superior robustness to model misspecification errors compared to the CG approach. The performance of the BIGPAST and CG methodologies is significantly influenced by the total variation distance between the skewed Student’s $t$ distribution and the normal distribution. When this distance is large, the BIGPAST approach exhibits superior accuracy compared to the CG method. The skewed Student’s $t$ distribution’s asymmetry, characterised by its light and heavy tails, further emphasises the necessity of employing a two-sided test over a one-sided test.
@@
The calculation of skewed Student’s $t$ distribution is implemented in the Python package skewt-scipy (https://pypi.org/project/skewt-scipy/). The code for simulation studies can be found in GitHub repository: BIGPAST. The real data comes from the Cambridge Centre for Ageing and Neuroscience (Cam-CAN) dataset (Shafto et al., 2014).
@@
The authors from the University of Kent, J.L., J.Z., and P.L. have no competing interests. G.G. and S.C. are employees of Innovision IP Ltd which provides commercial reports on individuals who may have had a head injury.

### v4

@@ 3.4 Comparison Study with Other Frameworks
This section is dedicated to assessing the performance of the proposed BIGPAST methodology and existing approaches when a control group is present. The first approach is a variant of BIGPAST based on the maximum a posteriori (MAP) estimation displayed in steps $3^{\prime}$ and $4^{\prime}$.
The second approach is a non-Bayesian approach that relies on maximum likelihood estimation (MLE). This approach assumes that the control group follows a skewed Student $t$ distribution. The MLE procedures can be outlined as follows:
@@
    Given the significance level $\beta$, the confidence interval of $x^{*}$ given $\mathbf{x}$ is $[x^{\mathrm{MLE}}_{\beta/2},x^{\mathrm{MLE}}_{1-\beta/2}]$, where $x^{\mathrm{MLE}}_{\beta/2}$ and $x^{\mathrm{MLE}}_{1-\beta/2}$ are the $\beta/2$ and $21-\beta/$ 2 quantiles of $f_{c}(\hat{\alpha}_{\mathrm{MLE}},\hat{\nu}_{\mathrm{MLE}},\hat{\xi}_{\mathrm{MLE}},\hat{\omega}_{\mathrm{MLE}})$, respectively. If $x^{*}$ does not fall within this confidence interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\beta$.
The third approach is a non-parametric (NP) approach. Given group data $\mathbf{x}$ and a significance level $\beta$, we can derive the empirical confidence interval $[x^{\mathrm{NP}}_{(\lfloor n*\beta/2\rfloor)},x^{\mathrm{NP}}_{(\lceil n*(1-\beta/2)\rceil)}]$ where $x^{\mathrm{NP}}_{(\lfloor n*\beta/2\rfloor)}$ and $x^{\mathrm{NP}}_{(\lceil n*(1-\beta/2)\rceil)}$ are the empirical quantiles of $\mathbf{x}$ at $\beta/2$ and $21-\beta/$ 2, respectively. If $x^{*}$ does not fall within this empirical confidence interval, then the null hypothesis $\mathcal{H}_{0}$ can be rejected at the significance level $\beta$. The NP approach does not rely on any distributional assumptions. Given that our BIGPAST methodology employs the Metropolis-Hastings sampler in Step 3, we will refer to the four approaches above: MH, MAP, MLE, and NP, respectively.
@@
$\mathrm{FDR}_{i}^{\mathcal{M}}\coloneqq\frac{\mathrm{FP}_{i}^{\mathcal{M}}}{\mathrm{TP}_{i}^{\mathcal{M}}+\mathrm{FP}_{i}^{\mathcal{M}}},\quad\mathrm{ACC}_{i}^{\mathcal{M}}\coloneqq\frac{\mathrm{TP}_{i}^{\mathcal{M}}+\mathrm{TN}_{i}^{\mathcal{M}}}{K},$
where $i=1,2,\ldots,N$, $\mathrm{TP}$, $\mathrm{FP}$, and $\mathrm{TN}$ represent the number of true positive, false positive and true negative, respectively. $\mathcal{M}$ could be one of MH, MAP, MLE, NP.
The parameters settings are as follows: we consider ten combinations of $(\alpha,\nu)$: (0,3), (1,3), (3,3), (5,5), (10,10), (20,20), (30,30), (50,50), (5,50) and (50,5) with fixed $\xi=-2$ and $\omega=2$. In MH approach, we let $m=5000$, $s=100$ and $c_{\nu}=1$. The other parameter settings are $N=400$, $n=100$, $k_{1}=k_{2}=1000$, $\beta=0.05$. The results are presented in Table 5 and Figure 5.
Table 5 displays the FDR and ACC of the four approaches. The MH and MAP methods consistently outperform the NP and MLE methods in terms of FDR and ACC. When ${\left\lvert\alpha\right\rvert}\leq 3$ and $\nu\leq 3$, both the FDR and ACC of the MAP approach surpass those of the MH approach. This is expected, as the MAP approach provides more accurate estimation of the underlying parameters of the skewed Student $t$ distribution when ${\left\lvert\alpha\right\rvert}$ and $\nu$ are small, as is shown in column 3 of Table 2. When ${\left\lvert\alpha\right\rvert}$ or $\nu$ is large, the FDR and ACC of the MH approach perform better than those of the MAP approach. This can be attributed to the randomness of parameters in the Metropolis-Hastings sampler. However, the MAP approach fails to estimate the underlying parameters when $\alpha$ or $\nu$ is large, as is shown in column 4 of Table 2, and lacks randomness of parameters. The boxplots in Figure 5 clearly illustrate the comparison between the MH and MAP approaches. The results demonstrate that the FDR of MH (i.e. BIGPAST) is generally lower than others, and the ACC of MH (i.e. BIGPAST) is generally greater than other methods in conducting hypothesis tests for single-subject data.
Table 5: The average false discovery rate (FDR) and accuracy (ACC) for the MH, MAP, MLE, and NP methods were calculated over 400 repetitions. In this context, MH refers to the original BIGPAST method, MAP denotes the Bayesian inference framework based on maximum a posteriori estimation, MLE represents the method based on maximum likelihood estimation, and NP signifies the non-parametric method.
@@ 4 Real Data Analysis
This dataset is derived from a mild traumatic brain injury (mTBI) study conducted by Innovision IP Ltd. It comprises a functional time series of Magnetoencephalography (MEG) data and an anatomical T1-weighted Magnetic Resonance Imaging (MRI) scan for 103 healthy subjects and one patient with mild traumatic brain injury. The MEG data, recorded from 306 sensors, was pre-processed using the ‘MNE-Python’ package (Gramfort et al., 2013). The MEG data was analysed to reconstruct the source generators using a minimum norm inverse method. The MRI data was pre-processed with the ‘FreeSurfer’ package (Han et al., 2006). The standard FreeSurfer pipeline (detailed on their website) was followed and includes motion correction, intensity normalisation, skull-stripping, normalisation to template space and segmentation of the cortex and subcortical structures using the Desikan-Killiany atlas. This results in 68 regions covering the entire cortex. For each cortical region, we calculate the average of source powers within this region. This process yields a control group dataset with dimensions $103\times 68\times 100$ and a patient dataset with dimensions $68\times 100$. To simplify, we treat the epochs as repeated observations and take averages over them to reduce the data’s complexity. Consequently, the control group dataset has dimensions $103\times 68$, and the patient dataset is reduced to a vector of length $68$.
@@
For each brain region of the Desikan-Killiany cortical atlas, we apply the BIGPAST method to the control group and patient dataset to test whether the patient significantly differs from the control group. This real data analysis shows that the single-subject has mTBI. However, in practice, we can calculate the observation of a new client and conduct the hypothesis test using BIGPAST to determine whether the client significantly differs from the control group.
We further explore the impact of waveband frequency on the comparison between a control group and a single-subject. The wavebands are arranged in ascending order of frequency: delta (1-4 Hz), theta (4-8 Hz), alpha (8-12 Hz), beta (12-30 Hz), and gamma (30-45 Hz). The subsequent results also include those of the Crawford-Garthwaite method, denoted as CG for brevity. The prespecified significance level is 0.05. The test results of Crawford-Garthwaite (CG) and BIGPAST are presented in Table 6, where the bracket (Normal, mTBI) in the third row represents that the result of CG is Normal, but the result of BIGPAST is mTBI. The numbers in cells are the indices of human brain regions; for example, number 1 in the “delta” column represents the caudal anterior cingulate in the left brain hemisphere. For this brain region, CG’s result is Normal, while the result of BIGPAST is mTBI. We did not consider the false discovery rate adjustment here as (1) steps 3 and 4 in BIGPAST employ the nested sampling technique, and the p-values are hard to obtain. (2) The aim of Table 6 is to provide a comparison between the CG and BIGPAST methods, and the FDR adjustment is not necessary for this purpose.
Table 6: The test results of the Crawford-Garthwaite (CG) based on Gaussian assumption and BIGPAST approaches based on the skewed Student $t$ assumption. The numbers in cells corresponding to specific brain cortical regions are 1: lh-caudal anterior cingulate, 2: lh-caudal middle frontal, 4: lh-entorhinal, 6: lh-fusiform, 10: lh-isthums cingulate, 12: lh-lateral orbitofrontal, 14: lh-medial orbitofrontal, 17: lh-parahippocampal, 22: lh-postcentral, 24: lh-precentral, 26: lh-rostral anterior cingulate, 35: rh-caudal anterior cingulate, 38: rh-entorhinal, 40: rh-fusiform, 47: rh-lingual, 50: rh-paracentral, 51: rh-parahippocampal, 58: rh-precentral, 59: rh-precuneus, 62: rh-superior frontal. The prefixes ‘lh-’ and ‘rh-’ indicate the left and right hemispheres of the brain, respectively. ‘NA’ stands for ‘not available’, and ‘Others’ represents the remaining brain cortical indices not listed in the column but can be found in the indices in Table 11.
The Results of (CG, BIGPAST) Wavebands
delta theta alpha beta gamma
@@
Figure 6: The comparison results between the Crawford-Garthwaite (CG) and BIGPAST approaches in terms of density estimations. The light-red lines represent the skewed Student $t$ densities, with parameters drawn from the BIGPAST approach’s posterior distribution. The light green lines depict the Normal densities, with parameters drawn from the CG approach’s posterior distribution. The CG test is based on 10,000 densities, while the BIGPAST test is based on 5,000 densities (the first 5,000 densities are discarded for burn-in). For clarity, each panel only displays 200 densities drawn from the CG and BIGPAST posterior distribution. The red vertical line on the $x$ axis represents the observation of a single-subject (i.e., a patient with mTBI). A histogram is used to demonstrate the goodness of fit for both the CG and BIGPAST approaches.
The null hypothesis in this real data analysis is two-sided, with one heavy tail and one light tail in the skewed Student $t$ distribution. The cortical region indices in row 3 of Table 6 exhibit a positive skewness parameter, i.e., $\alpha>0$. Consequently, the light tail is on the left, and the heavy tail is on the right. For all cortical region indices in row 3 of Table 6, the single-subject observation is located at the light tail of the skewed Student $t$ distribution. This positioning suggests a high probability of rejecting the null hypothesis, see Figure 6(a). Similarly, for cortical regions 6 (lh-fusion, gamma band), 14 (lh-medial orbitofrontal, theta and beta bands), 40 (rh-fusiform, theta and gamma bands), and 47 (rh-lingual, beta band), the single-subject observation is located at the heavy tail of the skewed Student $t$ distribution, as shown in Figure 6(b). The skewed Student $t$ assumption is more appropriate for these cortical regions than the assumption of Normal distribution. The BIGPAST approach outperforms the CG approach in detecting the single-subject observation at the heavy tail of the skewed Student $t$ distribution, as demonstrated in Figure 6. Because the BIGPAST approach corrects the model misspecification error at both tails compared to the CG approach. By combining rows 2 and 3 of Table 6 and Figure 6, we conclude that the BIGPAST approach is more reliable than the CG approach in hypothesis tests for single-subject data when the control group data exhibits skewness.
@@ 5 Conclusion
In this paper, we proposed a novel Bayesian inference framework for the abnormality detection on a single-subject versus a control group. Under the assumption of a Normal distribution, the $t$-score (Crawford and Howell, 1998) is capable of managing scenarios where the sample size is fewer than 30 (Machin et al., 2018). However, in neuroscience, the number of epochs often exceeds 100, sometimes even reaching 1000, and the distributions, whether over epochs or subjects, exhibit skewness. Under these conditions, methods predicated on the assumption of normality cease to be effective. Our proposed methodology, BIGPAST, leverages the skewed Student $t$ distribution, effectively capturing the inherent asymmetry of the data.
Our proposed BIGPAST methodology outperforms the existing Crawford-Garthwaite (CG) approach in terms of accuracy, power, and Type I error. Particularly when the control group data exhibits skewness, BIGPAST provides a more reliable solution for hypothesis testing on single-subject data. Furthermore, BIGPAST demonstrates superior robustness to model misspecification errors compared to the CG approach. The performance of the BIGPAST and CG methodologies is significantly influenced by the total variation distance between the skewed Student $t$ distribution and the Normal distribution. When this distance is large, the BIGPAST approach exhibits superior accuracy compared to the CG method. The skewed Student $t$ distribution’s asymmetry, characterised by its light and heavy tails, further emphasises the necessity of employing a two-sided test over a one-sided test.
@@
The calculation of skewed Student $t$ distribution is implemented in the Python package skewt-scipy. The code for simulation studies can be found in GitHub repository: BIGPAST. The real data comes from the Cambridge Centre for Ageing and Neuroscience (Cam-CAN) dataset (Shafto et al., 2014).
@@
The authors from the University of Kent, J.L., J.Z., and P.L. have no competing interests. G.G. and S.C. are employees of Innovision IP Ltd, which provides commercial reports on individuals who may have had a head injury.